OpenStudy (anonymous):

Pythagorean theorem question! A=? B=13 c=21

OpenStudy (anonymous):

a^2+b^2 = c^2 a^2 + 13^2 = 21^2 a^2 + 169 = 441 So 441 - 169 = 272 272 = a^2

OpenStudy (anonymous):

Thank youu

OpenStudy (anonymous):

your welcome : )